OpenStudy (anonymous):

how do you graph a line with a slope of 2: passing through (-3,0)

OpenStudy (anonymous):

put it in point slope form y-0=2(x-(-3) then it should be simplyfied to y-0=2(x+3) then convert to slope intercept y=2x+3 (because you distribute and add 0 to bth sides of the equation. then go to the point (0,3) and insert the slope (go up twice and to the right once)

OpenStudy (anonymous):

simply use the formula y-y1=m(x-x1) where m is the slope and (x1,y1) are the coordinates of the point

OpenStudy (anonymous):

@kmathsmiles55 how did you get the whole " y-0=2(x-(-3) " though?

OpenStudy (anonymous):

its an equation of a line. point slope form. it's y-y1=m(x-x1) the ones are little i cant find the button but its a real easy way to find what you're looking for becasue everything just kinda falls into place...you know what im saying?

OpenStudy (anonymous):

Ahhh. yeaah. I got it.

OpenStudy (anonymous):

WHOOOWWHH YAY!